OpenStudy (anonymous):

The circle intersects the line with he equation x+y=3 at 2 points, A and B. Find algebraically the coordinates A and B show the distance AB= sqrt 162 thanks :)

OpenStudy (anonymous):

what line?

OpenStudy (anonymous):

a circle equation has an equation of x6@+y^2=45 is that what you mean?

OpenStudy (anonymous):

MissMys, you just wrote something in the comments? is that more info about this problem?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Well you have to write the whole question word for word as i was given to you, in order for us to be of any help.

OpenStudy (anonymous):

a circle equation has an equation of x^2+y^2=45 The circle intersects the line with he equation x+y=3 at 2 points, A and B. Find algebraically the coordinates A and B show that the distance AB is sqrt 162 and the previous qs asks about radius and centre of the circle

OpenStudy (anonymous):

Put the two equations together, you would find to values for y=3, y=34. You can use these values to obtain related values for x. The two x,y values should be point A and B. You can double check it with distance formula.

OpenStudy (anonymous):

sorry but i dont understand? :S so do i put those two values into the equation??

OpenStudy (anonymous):

You have your equation for the line, you can solve for x, or y. From your line equation, x=3-y. Put this in the circle equation. (3-y)^2+y^2=45. I have already done this much for you, the result is y=3, y=34

OpenStudy (anonymous):

urm, i havent got a clue what you did and 34 is wrong :S i am so dead tomorrow :(

OpenStudy (anonymous):

i understand the first bit then confused of how you got 3 and 34 :S

OpenStudy (anonymous):

Well forget about 3 and 34. Solve for y from the new equation created.

OpenStudy (anonymous):

thanks for helping but i cant do it :( ill still give u the medal :)

OpenStudy (anonymous):

What can't you do? You have an equation (3-y)^2+y^2=45, solve for y.

OpenStudy (anonymous):

yea i dunno what or how you do the squared bit :(

OpenStudy (anonymous):

\[(3-y)^{2}=(3-y)(3-y)\]

OpenStudy (anonymous):

thanks :)